OpenStudy (anonymous):

Solve: 3e2x + 2 = 50 A. ln 8 B. ln 0.25 C. ln 14 D. 0.5ln 16

OpenStudy (anonymous):

algebra \[3e^{2x}+2=50\]\[3e^{2x}=48\]\[e^{2x}=16\]write in equivalent logarithmic form to get the variable out the exponent \[2x=\ln(16)=\ln(2^4)=4\ln(2)\]divide by 2\[x=2\ln(2)\]

OpenStudy (anonymous):

since this is not one of your "answer choices" i guess you need to stop at \[2x=\ln(16)\] and then divide by 2 to get \[x=.5\ln(16)\] which is a rather dumb way to write \(2\ln(2)\)

OpenStudy (anonymous):

Thanks so much!!